Please show me the solution to this problem using factoring and the zero-product property.
4a^2-10a +6=0
Thank you for helping me.

Answer by addingup(3677)

Factoring method:
4a^2-10a+6 = 0
We can factor the left side like this:
2(a-1)(2a-3) = 0
Now divide both sides by 2:
(a-1)(2a-3) = 0
a-1 = 0 or 2a-3 = 0
a = 1 or 2a = 3
a = 1 or a = 3/2